Solve. 3(x + 6) + 6x = 7x + 2
Answer by ankor@dixie-net.com(15652)

Here it is step-by-step
:
3(x + 6) + 6x = 7x + 2
:
Multiply what's inside the brackets:
3x + 18 + 6x = 7x + 2
:
Add the x's on the left
9x + 18 = 7x + 2
:
Subtract 18 from both sides
9x + 18 - 18 = 7x + 2 - 18
:
9x = 7x - 16
:
Subtract 7x from both sides:
9x - 7x = 7x - 7x - 16
:
2x = -16
:
1x = -16/2; divide both sides by 2
:
x = -8
:
:
Check our solution in the original equation substitute -8 for x:
3(x + 6) + 6x = 7x + 2
:
3(-8 + 6) + 6(-8) = 7(-8) + 2
:
3(-2) - 48 = -56 + 2
:
-6 - 48 = -56 + 2
:
-54 = -54, equality confirms our solution
:
Can you do these kind of equations now?